The following data show the probability distribution of the number of speakers sold daily in an appliance shop during the past several months, X={0, 1, 2, 3, 4, 5} P(X)={.07, .18, .27, .19, .08, .21}, what is the probability that at most 4 speakers will be sold?

1
Expert's answer
2022-03-07T10:21:04-0500

"P(X\u22644)=1-P(X>4)=1-P(X=5)=0.79"